OpenStudy (anonymous):

Kavita wants to borrow an amount of Rs.50000 for 3 years.Bank A charges an compund interest at 8% per annum ,while bank B charges simple interest at 5% semi anually .which offer is more profitable for Kavita.(show the whole solution)

OpenStudy (vishweshshrimali5):

ok let me solve this

OpenStudy (vishweshshrimali5):

see BANK A CI = 50000 (1 + 8/100)^ 3 = 62985.6 BANK B SI = 50000 * 10 * 3/ 100 = 15000 Clearly the BANK A will charge more. Thus, she should go for bank B

OpenStudy (vishweshshrimali5):

hope u got it now

OpenStudy (anonymous):

sorry ur sum is wrong because A = 50000 (1 + 8/100)^ 3 = 62985.6 AND CI = A - P = 12985.6 Therefore it is wrong